### v5

@@ 6.2. The nondeterminism hierarchy and GFMness
Let $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, be a finite memory strategy in ${\mathcal{F}}({\mathcal{M}})$, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain induced when taking actions in $\mathcal{M}$} according to $\mu$. Note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise. Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to $\mathcal{C}$}. Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of $\mathcal{N}$} over the word $\tau(r)$ that is sampled when playing in $\mathcal{M}$} with the strategy $\mu$. I.e., given $\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\ldots,\left\langle s_{k},m_{k}\right\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random. Notice that the composition $r\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},m_{0},q_{0}\right\rangle,\left\langle s_{1},m_{1},q_{1}\right\rangle,\ldots\right\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{N}}$ with a set of positions $S\times M\times Q_{\mathcal{N}}$. Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$ can be expressed as follows:
@@
We show that $\mathcal{N}$} is GFM. Consider an MDP $\mathcal{M}$}, and let $\mu:S\times M\to dist(A)\times M$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$. Consider the MC ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$, and note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a finite MC with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{a\in A_{s}}d(a)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$.
To conclude the GFMness of $\mathcal{N}$}, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$. Recall that a strategy for ${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in $\mathcal{M}$} and in $\mathcal{N}$} simultaneously. In the first it needs to choose an action $a\in A_{s}$ where $s\in S$ is the current state in $\mathcal{M}$}, while in the latter it should choose a $\tau(s)$-successor of the current state $q$ in $\mathcal{N}$}. The strategy $\mu^{\prime}$ resolves the first according to $\mu$, and hence ignores the $\mathcal{N}$} component. Then, for the $\mathcal{N}$} component, $\mu^{\prime}$ generates a run $r_{\mathcal{N}}$ of $\mathcal{N}$} on $w_{\mathcal{C}}$ as follows. The run $r_{\mathcal{N}}$ waits at $q_{0}$ (possibly forever), and moves to $q_{acc}$ only when an ergodic SCC that does not contain an $a$-labeled state is reached in $\mathcal{C}$}; in particular, if the ergodic SCC that is reached in $\mathcal{C}$} contains an $a$-labeled state, then $r_{\mathcal{N}}$ does not leave $q_{0}$. In other words, the strategy $\mu^{\prime}$ moves from $q_{0}$ to $q_{acc}$ only when it is certain that there are no more $a$’s in the suffix to be read.
@@ Appendix B Details of the Reduction in Theorem 3
  • $next(\#,(q,\gamma_{i}),\gamma_{i+1})=(q^{\prime},\gamma^{\prime})$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},L)$ 111111 We assume that the reading head of $T$ stays in place when it is above the leftmost cell and the transition function directs it to move left..

### v6

@@ 6.2. The nondeterminism hierarchy and GFMness
Let $\mu:S\times M\rightarrow\mathrm{dist}(A)\times M$, be a finite memory strategy in ${\mathcal{F}}({\mathcal{M}})$, and let ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$ be the Markov chain induced when taking actions in $\mathcal{M}$} according to $\mu$. Note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a Markov chain with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined by $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{c\in A_{s}}d(c)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,c,s^{\prime})$, when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$, otherwise. Let $r=\left\langle\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\left\langle s_{2},m_{2}\right\rangle,\ldots\right\rangle\in(S\times M)^{\omega}$ be a random process sampled according to $\mathcal{C}$}. Let $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{\omega}$ be a random run of $\mathcal{N}$} over the word $\tau(r)$ that is sampled when playing in $\mathcal{M}$} with the strategy $\mu$. I.e., given $\left\langle s_{0},m_{0}\right\rangle,\left\langle s_{1},m_{1}\right\rangle,\ldots,\left\langle s_{k},m_{k}\right\rangle$, for $k\geq 1$, and $q_{0},q_{1},\ldots,q_{k-1}$, we set $q_{k}$ to be one of the $\tau(s_{k})$-successors of $q_{k-1}$ at random. Notice that the composition $r\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},m_{0},q_{0}\right\rangle,\left\langle s_{1},m_{1},q_{1}\right\rangle,\ldots\right\rangle$ is a process of a finite Markov chain ${\mathcal{C}}_{\mathcal{N}}$ with a set of positions $S\times M\times Q_{\mathcal{N}}$. Indeed, the stochastic transition function of ${\mathcal{C}}_{\mathcal{N}}$ can be expressed as follows:
@@
We show that $\mathcal{N}$} is GFM. Consider an MDP $\mathcal{M}$}, and let $\mu:S\times M\to dist(A)\times M$ be a finite memory strategy such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$. Consider the MC ${\mathcal{C}}={\mathcal{C}}({\mathcal{M}},\mu)$, and note that since $\mu$ uses $M$ as a memory structure, we may think of $\mathcal{C}$} as a finite MC with positions $S\times M$, and a stochastic transition function $\mathrm{Pr}_{\mathcal{C}}:(S\times M)^{2}\rightarrow[0,1]$ that is defined $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=\sum_{c\in A_{s}}d(c)\cdot\mathrm{Pr}_{{\mathcal{M}}}(s,c,s^{\prime})$ when $\mu(s,m)=\left\langle d,m^{\prime}\right\rangle$, and otherwise $\mathrm{Pr}_{\mathcal{C}}(\left\langle s,m\right\rangle,\left\langle s^{\prime},m^{\prime}\right\rangle)=0$.
To conclude the GFMness of $\mathcal{N}$}, we show that there is a strategy $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$. Recall that a strategy for ${\mathcal{M}}\times{\mathcal{N}}$ needs to take actions in $\mathcal{M}$} and in $\mathcal{N}$} simultaneously. In the first it needs to choose an action $c\in A_{s}$ where $s\in S$ is the current state in $\mathcal{M}$}, while in the latter it should choose a $\tau(s)$-successor of the current state $q$ in $\mathcal{N}$}. The strategy $\mu^{\prime}$ resolves the first according to $\mu$, and hence ignores the $\mathcal{N}$} component. Then, for the $\mathcal{N}$} component, $\mu^{\prime}$ generates a run $r_{\mathcal{N}}$ of $\mathcal{N}$} on $w_{\mathcal{C}}$ as follows. The run $r_{\mathcal{N}}$ waits at $q_{0}$ (possibly forever), and moves to $q_{acc}$ only when an ergodic SCC that does not contain an $a$-labeled state is reached in $\mathcal{C}$}; in particular, if the ergodic SCC that is reached in $\mathcal{C}$} contains an $a$-labeled state, then $r_{\mathcal{N}}$ does not leave $q_{0}$. In other words, the strategy $\mu^{\prime}$ moves from $q_{0}$ to $q_{acc}$ only when it is certain that there are no more $a$’s in the suffix to be read.
@@ Appendix B Details of the Reduction in Theorem 3
  • $next(\#,(q,\gamma_{i}),\gamma_{i+1})=(q^{\prime},\gamma^{\prime})$ where $(q,\gamma_{i})\rightarrow(q^{\prime},\gamma^{\prime}_{i},L)$  [^11] We assume that the reading head of $T$ stays in place when it is above the leftmost cell and the transition function directs it to move left..